Find the mean and the Variance of the following sample data:             x Frequency (f) 1...

  1. Find the mean and the Variance of the following sample data:

           

x

Frequency (f)

1

5

2

6

4

9

8

6

12

4
Answer #1
x Frequency (f) Relative frequency x^2 x(px) x^2(pX)
1 5 0.1667 1 0.166666667 0.166666667
2 6 0.2 4 0.4 0.8
4 9 0.3 16 1.2 4.8
8 6 0.2 64 1.6 12.8
12 4 0.1333 144 1.6 19.2
sum= 37.7667
Mean= 4.9667

Variance =

V(x) = E(x^2) - E(x)^2 =
37.7667 - 24.6681 = 13.0986